OpenStudy (anonymous):

If possible, solve each system of equations. Use any method. If there is not a unique solution to a system, state a reason. 2x-8y=5 -6+24y=15

OpenStudy (anonymous):

double check and make sure your second equation is typed properly.

OpenStudy (anonymous):

if it is correct you simply solve the bottom equation for y and plug that value into the first equation's "y"

OpenStudy (anonymous):

my problem is right now i need help please

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

add 6 to both sides and divide both sides by 24, that will isolate 'y'

OpenStudy (anonymous):

I don't see it can you please show me

OpenStudy (anonymous):

\[-6+24y=15\] add 6 to both sides \[24y = 21\] divide both sides by 24\[y = \frac{ 21 }{ 24}\] which reduces to \[y = \frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

now you take that value and plug it into your other equation, like so: \[2x-8*(\frac{ 7 }{ 8})=5\]

OpenStudy (anonymous):

can you take it from there?

OpenStudy (anonymous):

i got it. thank you

OpenStudy (anonymous):

no problem, good luck!

OpenStudy (anonymous):

thank you